OpenStudy (anonymous):

Rachel moved into a new apartment m months ago and paid a $1415 deposit. She pays a monthly rent of $750. How many months has she lived in her apartment if she has paid $8,165? Numerical Answers Expected!

OpenStudy (mathteacher1729):

@Snapbacklive How far have you gotten with this problem so far? Where are you getting stuck with it?

OpenStudy (anonymous):

We can set this up as an equation. Let T be the total cost, and M be the number of months she's lived there. Then T = 750M + 1415. We have that she's paid $8165 so far, so plug that in for T and solve for M.

OpenStudy (anonymous):

so T= 8,165 + 750M

OpenStudy (mathteacher1729):

@Snapbacklive T is the total cost paid after M months. You're trying to find out: "How many months has she lived in her apartment if she has paid $8,165? " In other words, solve 8,165 = 750*M + 1415 for "M". Do you understand why this is the desired equation? (It's ok to be confused, no worries).

OpenStudy (anonymous):

So 10 ?

OpenStudy (anonymous):

@mathteacher1729 , @mslano

OpenStudy (mathteacher1729):

@Snapbacklive how did you obtain the value of 10? Did you graph it? Did you guess & check? Did you solve it algebraically?

OpenStudy (anonymous):

i used a calculator

OpenStudy (anonymous):

can u please me on this other question please.

OpenStudy (mathteacher1729):

What does 10 represent? Also what specific steps did you take on your calculator? It's difficult to give help if I don't know what you're doing, mathematically.

OpenStudy (anonymous):

i just put 750 times 10.888 but it was wrong

OpenStudy (mathteacher1729):

Here's a step-by-step on how to solve 8,165 = 750*M + 1415 for "M" Step one: 8,165 = 750*M + 1415 now subtract 1415 from both sides and what do you get?

OpenStudy (anonymous):

it's about 11 months

OpenStudy (mathteacher1729):

Depends on how you define "about". I mean, 11 months is closer than say 140 months... ;-p But ok. 8,165 = 750*M + 1415 subtract 1415 from both sides... 6,750 = 750*M now divide both sides by 750

OpenStudy (anonymous):

now can u please help me on this other question please.

OpenStudy (anonymous):

@mathteacher1729

OpenStudy (mathteacher1729):

@Snapbacklive ... I like to focus on one thing at a time, sorry.

OpenStudy (anonymous):

ok